# Universal on-chip polarization handling with deep photonic networks

### v2

@@ Abstract
In recent years, guided-wave platforms, particularly silicon-on-insulator photonics, have experienced significant advancement with applications expanding from traditional communications and sensing to next-generation spectroscopy, optical computation, and artificial intelligence [1, 2, 3, 4, 5, 6, 7]. In these high-data-throughput applications, managing optical inputs with varying polarizations continues to be a universal and key requirement [8, 9, 10, 11]. Inputs with different polarizations experience different responses on-chip due to the high-index contrast and specific waveguide aspect ratios, leading to performance degradation in the form of polarization-dependent losses as well as changes in waveguide dispersion and spectral characteristics. As these performance limitations can be detrimental to large-scale photonic systems, effective polarization management remains an essential requirement for most on-chip photonic applications [12].
@@
Previous studies have explored handling of different polarizations through a number of methods, yielding devices that are either polarization-independent or capable of selectively managing signals based on their polarization state [13], based on asymmetrical couplers [14, 15], microring resonators [16], bent directional couplers [17] and interferometers [18, 19]. Examples include polarization-independent power splitters using mode evolution [20], polarization splitters and rotators with subwavelength grating claddings [21], and polarization splitters achieved through adiabatic couplers [22], multi-level tapers, or counter-tapered couplers [23, 24]. Polarization demultiplexing has also been demonstrated using integrated dielectric metasurfaces [25]. As these prior examples are designed for application-specific tasks, they are typically limited in their functional flexibility and operation bandwidths since they rely on specific waveguide geometries utilizing wave coupling dynamics. Furthermore, while these designs address particular polarization handling tasks, the diversity in their waveguide architectures and application-specific parameters points to the lack of a universal design framework for handling different input polarization states. Such a framework would ideally be adaptable to a wide range of functionalities and material platforms, enabling arbitrary polarization handling capabilities on-chip.
@@
In this paper, we introduce a universal design paradigm for arbitrary polarization management using silicon-based deep photonic networks of cascaded Mach-Zehnder interferometers (MZIs), capable of handling both TE and TM polarizations over a wide range of wavelengths on the same platform. Each MZI incorporates custom tapers with specific phase profiles that collectively build the transfer function of the entire network based on input polarization and wavelength. Leveraging a physics-informed neural framework and a gradient-based optimization algorithm, we iteratively train these networks to achieve user-specified optical responses. We demonstrate this capability with three broadband devices: a polarization splitter, a polarization-independent power splitter, and a polarization-dependent arbitrary coupler. With design times under one minute per device, these photonic networks exhibit state-of-the-art performance, including flat-top transmission bands, extinction ratios of 15–20 dB, and operational bandwidths exceeding 120 nm, as verified by 3D-FDTD simulations. Collectively, these devices provide a broad range of polarization management functionalities, highlighting the adaptability of our photonic network design within a unified system architecture. These capabilities can be extended to various material and fabrication platforms, positioning deep photonic networks as an effective solution for enhancing throughput in next-generation optical applications requiring versatile polarization control.
@@ II-B Optimization Flow
The optimization procedure for designing universal polarization handling deep photonic networks is illustrated in Fig 2 with two main sequential steps: device initialization and optimization loop. First, the deep photonic network architecture including the number of input/output ports, the network depth (number of layers and MZIs), and the optical objectives of the network are defined for TE and TM polarizations and multiple wavelengths. Subsequently, an artificial, software-defined equivalent of the deep photonic network is created using automatic differentiation-compatible building blocks as described above, the architectural hyperparameters from the preceding phase, and an end-to-end machine learning library [30].
@@
The gradient of this L2-loss is calculated using backpropagation through the network layers and then used to iteratively update the trainable waveguide widths to minimize the difference between the target optical functionality and the simulated device response. This systematic solution for the optimal device geometry is completed once the device achieves a near-perfect match to the target response at both polarizations, depending on the specified convergence criteria. During this solution, several regularizations of the resulting waveguide widths can be used in order to limit potential losses in the custom tapers. To minimize potential scattering losses due to abrupt width transitions, the widths are constrained between 350 nm and 700 nm, around a standard width of 450 nm. Furthermore, several regularizations applied during the design process collectively reduce abrupt changes in waveguide and guide them toward the 450 nm default, ensuring a reliable and low-loss final device design. For the specific devices in this paper, these regularizations resulted in waveguide widths between 400 nm and 580 nm in the custom tapers. Following the optimization loop, the physical device layout is generated from the geometrical description, from which further verification through separate electromagnetic simulations or analyses of fabrication variations can be performed.
@@ III Polarization-Handling Device Designs and Simulations
Using the methodology outlined above, we demonstrate three distinct deep photonic networks as polarization handling devices, all in the form of 1-input 2-output splitters. All three networks were constructed using five layers of cascaded MZIs, resulting in a total footprint of for each device. To demonstrate the versatility of our design approach and showcase polarization-dependent capabilities, increasingly complex polarization-handling tasks were chosen for the three devices. Each device was designed for wideband operation for a bandwidth of 120 nm between 1490 nm and 1610 nm. The first device was designed as a polarization splitter where TE and TM inputs were separated from one another, and individually directed to their respective output ports. The second device was designed to evenly separate the input power into the two output ports regardless of polarization, acting as a polarization-independent 50/50 power splitter. The target operation for the last device was chosen as equal power distribution to the two outputs for only the TE-polarized input, while the TM-polarized input is directed to only one of the outputs. For each one of the devices, the transmission characteristics were compared with 3D-FDTD simulations; and fabrication tolerance analyses were performed. These devices and their analyses are detailed in the following sections.
@@ III-A Polarization Splitter
Next, we calculate the tolerance of the network to fabrication-induced variations and plot the transmission responses under waveguide width variations of up to 10 nm with shaded regions in Fig 3d. The nominal transmissions under ideal fabrication conditions are also plotted with the solid curves for reference. The results indicate that the extinction ratio remains over 15 dB and it is consistent with measurement results of other polarization splitters demonstrated previously [31, 32]. The results also show that the TM polarization response exhibits better fabrication tolerance compared to the TE polarization response, attributed to the slower rate of change in the effective index of TM polarization, which reduces its phase delay sensitivity to waveguide width changes.
@@ III-C Polarization-dependent Arbitrary Power Splitter
For more complex device objectives like this one, the deep photonic network design procedure enables the utilization of a greater number of layers, offering increased design flexibility. Even though five layers were sufficient for all objectives here, designers may easily experiment with the device performance and fabrication tolerance results in deeper networks with greater number of layers. The networks also generalize well to more than two outputs, where an arbitrary number of input-output pairs can be specified, each with its own transfer function. The simulation results obtained from these three devices underscore the efficacy of deep photonic networks in achieving arbitrarily specified polarization handling capabilities characterized by high extinction ratios and minimal insertion losses. We also note that these networks may be configured in order to reduce transmission errors due to fabrication-induced changes, by simultaneously optimizing over-etched and under-etched versions of the same device [27]. Moreover, the demonstrated methodology is not restricted to silicon and can be extended to other material platforms, depending on system requirements. By offering a framework that can be standardized and applied across different materials or fabrication platforms, this design paradigm enables universal polarization handling capabilities within a unified device architecture. These capabilities underscore deep photonic networks as a scalable and universally applicable architecture for realizing application-specific polarization handling operations across various domains including next-generation optical communications, sensing, and computing applications.
@@ IV Conclusion
In summary, we have presented a pioneering methodology for the design of photonic devices capable of arbitrary on-chip polarization manipulation through deep photonic networks. By harnessing custom tapers and a software-defined neural network framework, we presented the design of broadband devices with state-of-the-art performance metrics including flat-top transmission bands, high extinction ratios, and wide operational bandwidths. Specifically, our numerical results demonstrate extinction ratios of over 20 dB, and operational bandwidths exceeding 120 nm centered around 1550 nm across various devices for both polarization-independent and polarization-specific photonic tasks. The demonstrated scalability and adaptability of this design framework can advance integrated photonics capabilities, with potential applications in optical communication, sensing, and computing.
@@ Acknowledgments
This work was supported by the Scientific and Technological Research Council of Turkey (TUBITAK) under grant number 123E567, and the Marie Sklodowska-Curie Fellowship (no 101032147) through the Horizon 2020 program of the European Commission.

### v3

@@ Abstract
In recent years, guided-wave platforms, particularly silicon-on-insulator photonics, have experienced significant advancement with applications expanding from traditional communications and sensing to next-generation spectroscopy, optical computation, and artificial intelligence [1, 2, 3, 4, 5, 6, 7]. Managing varying input polarizations in these high-throughput applications remains a key requirement [8, 9, 10, 11]. Unintended polarization-dependent on-chip responses due to high-index contrast and waveguide aspect ratios lead to performance degradation including polarization-dependent losses and modified spectral characteristics. Since such limitations impair large-scale photonic systems, effective on-chip polarization management is essential for most applications [12].
@@
Previous studies have explored handling of different polarizations through a number of methods, yielding devices that are either polarization-independent or capable of selectively managing signals based on their polarization state [13], based on asymmetrical couplers [14, 15], microring resonators [16], bent directional couplers [17] and interferometers [18, 19]. Examples include polarization-independent power splitters using mode evolution [20], polarization splitters and rotators with subwavelength grating claddings [21], and polarization splitters achieved through adiabatic couplers [22], multi-level tapers, or counter-tapered couplers [23, 24]. Polarization demultiplexing has also been demonstrated using integrated dielectric metasurfaces [25]. These application-specific prior examples are often limited in flexibility and bandwidth due to their reliance on specific waveguide geometries and wave coupling dynamics. Moreover, the diverse architectures and parameters in these designs highlight the lack of a universal framework for managing different input polarizations. Such a framework would ideally be adaptable to a wide range of functionalities and material platforms, enabling arbitrary polarization handling capabilities on-chip.
@@
In this paper, we introduce a universal design paradigm for arbitrary polarization management using silicon-based deep photonic networks of cascaded Mach-Zehnder interferometers (MZIs), capable of handling both TE and TM polarizations over a wide range of wavelengths. Custom tapers with specific phase profiles in each MZI collectively form the network’s transfer function based on the input polarization and wavelength. Using a physics-informed neural framework and gradient-based optimization, we iteratively train these networks for user-specified optical responses. We demonstrate this capability with three broadband devices: a polarization splitter, a polarization-independent power splitter, and a polarization-dependent arbitrary coupler. Designed in under a minute per device, these networks show state-of-the-art performance with flat-top transmission, 15–20 dB extinction ratios, nm bandwidths, verified by 3D-FDTD simulations. Collectively, these devices provide a broad range of polarization management functionalities, highlighting the adaptability of our photonic network design within a unified system architecture. These capabilities can be extended to various material and fabrication platforms, positioning deep photonic networks as an effective solution for enhancing throughput in next-generation optical applications requiring versatile polarization control.
@@ II-B Optimization Flow
The optimization procedure for designing these deep photonic networks is illustrated in Fig. 2 with two main steps: device initialization and an optimization loop. First, the network architecture (input/output ports, numbers of MZI layers) and optical objectives are defined for TE/TM polarizations across multiple wavelengths. Then, a software-defined equivalent of this network is created using previously described automatic differentiation-compatible building blocks, architectural hyperparameters, and an end-to-end machine learning library [30].
@@
The gradient of this L2-loss is calculated via backpropagation, and used to iteratively update the trainable waveguide widths to minimize deviation from the target optical functionality. Optimal device geometry is achieved when the simulated response closely matches the target for both polarizations, per convergence criteria. During this optimization, several regularizations on waveguide widths are used to limit potential losses in custom tapers. To minimize potential scattering losses due to abrupt width transitions, the widths are constrained between 350 nm and 700 nm, around a standard width of 450 nm. Furthermore, several regularizations applied during the design process collectively reduce abrupt changes in waveguide and guide them toward the 450 nm default, ensuring a reliable and low-loss final device design. These regularizations yielded custom taper widths between 400 nm and 580 nm for our devices. Post-optimization, the physical layout is generated from its geometrical description for further simulation/verification and fabrication variation analysis.
@@ III Polarization-Handling Device Designs and Simulations
Using our methodology, we demonstrate three 1-input, 2-output deep photonic networks for polarization handling. To showcase design versatility, we design three devices with increasingly complex polarization-handling tasks over a 120 nm bandwidth (1490-1610 nm). All three networks are constructed from five MZI layers, resulting in a total footprint of for each device. The first device, a polarization splitter, separates TE and TM inputs, directing each to its respective output port. The second, a polarization-independent 50/50 power splitter, evenly divides the input power between the two outputs regardless of polarization. The final device equally splits TE-polarized input power to two outputs, while directing all of the TM-polarized input to only one of the output ports. For each device the transmission characteristics were verified with 3D-FDTD simulations; and fabrication tolerance analyses were performed.
@@ III-A Polarization Splitter
Next, we calculate the tolerance of the network to fabrication-induced variations and plot the transmission responses under waveguide width variations of up to 10 nm with shaded regions in Fig 3d. The nominal transmissions under ideal fabrication conditions are also plotted with the solid curves for reference. The results indicate that the extinction ratio remains over 20 dB and it is consistent with measurement results of other polarization splitters demonstrated previously [31, 32]. The results also show that the TM polarization response exhibits better fabrication tolerance compared to the TE polarization response, attributed to the slower rate of change in the effective index of TM polarization, which reduces its phase delay sensitivity to waveguide width changes.
@@ III-C Polarization-dependent Arbitrary Power Splitter
For more complex objectives, our design procedure allows for more cascaded layers with increasing design flexibility. While five layers were sufficient for all devices here, deeper networks can be easily explored for performance and fabrication tolerance. The networks also generalize to more than two outputs, with specifiable transfer functions for arbitrary input-output pairs. The simulation results from the three devices we study here underscore our networks’ capabilities in achieving arbitrarily specified polarization handling with high extinction ratios and minimal insertion losses. We also note that the same optimization procedure may also be used to mitigate fabrication-induced errors by simultaneously optimizing over/under-etched versions of the same device [27]. Moreover, the demonstrated methodology is not restricted to silicon and can be extended to other material platforms, depending on system requirements. By offering a framework that can be standardized and applied across different materials or fabrication platforms, this design paradigm enables universal polarization handling capabilities within a unified device architecture. These capabilities highlight deep photonic networks as a scalable, universally applicable architecture for specific polarization handling in advanced optical communications, sensing, and computing.
As detailed in Table I, our deep photonic network devices demonstrate competitive performance compared to state-of-the-art designs [13, 15, 31, 32, 37, 38, 20, 34, 35, 36]. Our polarization splitters achieve an extinction ratio exceeding 20 dB with less than 0.45 dB excess loss over a 120 nm bandwidth. The polarization-independent power splitter maintains a splitting imbalance below 1 dB and an excess loss under 0.5 dB across the same spectrum. Beyond their strong performance, our devices are also designed in under one minute each, demonstrating several orders of magnitude better computational efficiency than the designs for comparable structures. This efficiency offers a major advantage over conventional approaches, which require prior knowledge of mode coupling dynamics, and over other inverse design methods that rely on iterative electromagnetic simulations. The rapid optimization and flexibility of our method enable the design of complex, general-purpose polarization handling devices that are otherwise challenging to realize efficiently.
| Reference | Structure | Design Timeframe | Capability | Footprint [µm2] | ER [dB] | Split Imb. [dB] | EL [dB] | Bandwidth [nm] |
| Rotating/Non-Rotating Polarization Splitters | ||||||||
| [13] | Double-etched Taper | traditional design (hours) | function-specific | — | ||||
| [32] | Double-etched DC | traditional design (hours) | function-specific | — | 30 | |||
| [15] | Taper + ADC | traditional design (hours) | function-specific | — | ||||
| [31] | Multi-mode WG + ADC | traditional design (hours) | function-specific | — | ||||
| [37] | Inverse Design | 12 hr | general purpose | — | ||||
| [38] | Inverse Design | 19 hr | general purpose | — | ||||
| Our Work | Deep Photonic Network | 58.1 sec | general purpose | 1280 | — | 120 | ||
| Polarization-Independent Power Splitters | ||||||||
| [20] | Slotted WG + Tapers | traditional design (hours) | function-specific | — | ||||
| [34] | Modified MMI Junction | traditional design (hours) | function-specific | — | ||||
| [35] | SWGs | traditional design (hours) | function-specific | — | ||||
| [36] | Cascaded Bent DCs | traditional design (hours) | function-specific | — | ||||
| Our Work | Deep Photonic Network | 54.9 sec | general purpose | 1280 | — | 120 | ||
| ER: Extinction Ratio. Split Imb. (Splitting Imbalance): Power difference between the different output ports relative to input power. EL: Excess Loss. Bandwidth: 0.5 dB bandwidth, defined as the wavelength range where transmissions remain within 0.5 dB of their optimal values. ADC: Asymmetric Directional Coupler. MMI: Multi-Mode Interference. SWG: Subwavelength Grating. WG: Waveguide. | ||||||||
@@ IV Conclusion
In summary, we presented a pioneering deep photonic network methodology for designing devices for arbitrary on-chip polarization manipulation. Using custom tapers and a software-defined neural network framework, we designed broadband devices with state-of-the-art metrics including flat-top transmission, high extinction ratios, and wide operational bandwidths. Our numerical results specifically show ¿20 dB extinction ratios and ¿120 nm operational bandwidths (centered at 1550 nm) for various polarization-handling tasks. This design framework’s demonstrated scalability and adaptability can advance integrated photonics, with potential applications in optical communication, sensing, and computing.
@@ Acknowledgments
This work was supported by the Scientific and Technological Research Council of Turkey (TUBITAK) under grant number 119E195, and the Marie Sklodowska-Curie Fellowship (no 101032147) through the Horizon 2020 program of the European Commission.